An acorn falls from a tree that is 75 feet tall. What is the height of the acorn 1 second after it was dropped? Use the vertical motion model, h=-16t^2 + 75 where t is the time in seconds and h is the height in feet.
Height:_ feet

Answer by ikleyn(52780)

.

To answer this question, you should substitute t= 1 second into the formula and calculate the 


      height = -16*1^2 + 75 = -16 + 75 = 59 feet from the ground.    ANSWER

Solved.